WHICH RELATION IS NOT A FUNCTION
Y=X+2
Y=X^2-4
X=Y^2-4Y
X=3Y+4

Answer by stanbon(75887)

WHICH RELATION IS NOT A FUNCTION
Y=X+2
Y=X^2-4
X=Y^2-4Y
X=3Y+4
---------
x=y^2-4y is not a function
Why?
Let x=0
Then y^2-4y=0
y(y-4)=0
y=0 or y=4
With a single value of x there are two values of y.
Therefore the rule is not a function.
Cheers,
Stan H.